Chastain Corporation is trying to determine the effect of its inventory turnover ratio and days sales outstanding (DSO) on its cash conversion cycle. Chastain's 2019 sales (all on credit) were $197,000, its cost of goods sold is 80% of sales, and it earned a net profit of 4%, or $7,880. It turned over its inventory 7 times during the year, and its DSO was 33 days. The firm had fixed assets totaling $31,000. Chastain's payables deferral period is 50 days. Assume 365 days in year for your calculations.

  1. Calculate Chastain's cash conversion cycle. Do not round intermediate calculations. Round your answer to two decimal places.
    days

  2. Assuming Chastain holds negligible amounts of cash and marketable securities, calculate its total assets turnover and ROA. Do not round intermediate calculations. Round your answers to two decimal places.
    Total assets turnover:   
    ROA: %

  3. Suppose Chastain's managers believe that the inventory turnover can be raised to 8.6 times. What would Chastain's cash conversion cycle, total assets turnover, and ROA have been if the inventory turnover had been 8.6 for 2019? Do not round intermediate calculations. Round your answers to two decimal places.
    Cash conversion cycle: days
    Total assets turnover:   
    ROA:
Answer #1

a.Cash Conversion Cycle = DIO + DSO - DPO

DIO = ((Average Inventory)/ Cost of Sales) x 365

Average Inventory = 1/7 times the turnover = 197000 x 0.8.x (1/7) = 22,514.28

Cost of sales = 157600

DIO = (22,514.28/1,57,600) x 365 = 52.14

DSO = 33

DPO = 50 days

Cash Conversion cycle = 52.14 + 32 - 50 = 34.14 days

c.Asset Turnover Ratio = Sales / Assets = 197000/ 31000 = 6.35

RoA = Net Profit / Assets = 7880/31000 = 0.25

e. Revised Cash Conversion cycle based on 8.6 inventory turnover

DIO = Average Inventory / Cost of sales x 365

Average Inventory will be 197000/8.6 = 22906.97

DIO = 22906.97/157600 x 365 = 53.05 days

Cash conversion cycle =53.05 +32 - 50 = 35.05

Total Asset Turnover and RoA will remain the same.